OpenStudy (anonymous):

Which set of numbers could represent the lengths of the sides of a right triangle? Answer 3, 6, 9 4, 8, 12 3, 4, 5 3, 5, 8

OpenStudy (anonymous):

right angle triangle means one angle of 90 degree or follow the Pythagoras theorm so use law of cosine to solve.

OpenStudy (anonymous):

im still really confused

OpenStudy (anonymous):

3,4,5 are the sides of triangle

OpenStudy (anonymous):

@best.shakir why do you post answers only?

OpenStudy (anonymous):

3x+4x+5x=180 12x=180 x=180/12=15

OpenStudy (anonymous):

3*15+4*15+5*15=180